The length of a rectangle is 15 m more than its width. If the perimeter of rectangle is 50 m, find the dimensions of the rectangle.
Found 2 solutions by advanced_Learner, addingup:
Answer by advanced_Learner(501)

The length of a rectangle is 15 m more than its width. If the perimeter of rectangle is 50 m, find the dimensions of the rectangle.


50=2l+2w
50=2(w+15)+2w
5=(w)
20=(l)

Answer by addingup(3677)

Perimeter = 2L+2W
and the problem says:
L = W+15 Substitute above for L:
2(W+15)+2W = 50
2W+30+2W = 50 add W on left and subtract 30 on bboth sides
4W = 20
W = 5 The width is 5, and the length
L= 5+15 = 20
--------------------
Check:
2(20)+2(5) = 50
40+10 = 50 Correct
:
John